# Image Classification with Rotation-Invariant Variational Quantum Circuits

### v1

@@ III.5 Scaling the problem size
The data set employed in the previous sections represents a toy case created for the comparison of the capabilities of the geometric approach with respect to the traditional variational quantum models. The resolution of images in real computer vision problems, however, is far from the phew pixels the current NISQ-era devices can handle due to their limited circuit size. A way to address this issue is to use a another encoding strategy, different from the one proposed in this work, that optimizes the number of features encoded per qubit. In [44, 43], for example, Amplitude Encoding is used the encode the images into the quantum circuits. Another way to cope with this matter, rather than waiting to larger-scale quantum devices to be built and implement the proposed architecture for larger images, is to reduce the dimensionality of the input.
@@
where $r=(n_{f}-1)/2$ for odd $n_{f}$ and $r=n_{f}/2$ for even $n_{f}$, being $n_{f}$ the size of the filter. This expression can be extended to consider RGB channels, but in general, the operation is not equivariant to image rotations described by eq. 9. For this reason, following the approach of [57], a rotation-equivariant convolution operation is defined by computing a convolution between an image and all rotated versions of the original filter, and pooling all results (see fig. 9).
@@
These equivariant convolutional layers can be stacked arbitrarily in $n_{c}$ channels, each one of them with a different filter so that they learn to focus on specific characteristics of the images. This way, if $n_{x}$ is the size (number of pixels) of the rows of the square image, $n_{w}$ the size of the filter, $s$ the stride and $p$ the padding, the size of the convoluted image will be [58]:
@@
For this case, $p=0$ and $s=1$ will be taken, so $n^{\prime}_{x}=n_{x}-n_{w}+1$. Taking the data set as a tensor of $N$ data-points allows us to apply the equivariant convolution operation over the whole batch, performing the following dimensional reduction:
@@
Accordingly, the number of channels of each convolutional layer and the size of the filters can be tuned to reduce the number of features and adapt the input to the size of the available quantum circuit. In this case, to connect the classical part to the equivariant quantum classifier described in the previous sections, the original images will be reduced to $4\times 4$, single channel square images. It is postulated that while the number of qubits keeps increasing and the connectivity between them condensing, the size of the images that can be embedded directly into the quantum classifier will increase while decreasing the dependency on the classical convolutional part.
@@
This algorithm was tested on two public datasets. The first one, Fashion MNIST [^1] https://huggingface.co/datasets/fashion_mnist, contains $28\times 28$ grayscale images. For binary classification, labels 00 (T-shirts) and $1$ (trousers) have been taken and for a dataset of 200 images, convolution filters of dimensions of $n_{w}=\{11,11,3,3\}$ and $n_{l}=2$ data re-uploading layers, a mean performance of $F1_{\text{train}}=0.82\pm 0.09$ and $F1_{\text{test}}=0.79\pm 0.11$ was obtained. The second dataset, 11k Hands [^2] https://sites.google.com/view/11khands [59], has been chosen as an illustrative example of a rotational symmetric problem where reflection symmetry is not present. It contains $1600\times 1200$ RGB images of hands that have been reduced to a resolution of $64\times 64$ before applying the hybrid algorithm. 500 images of left and right hands have been taken for binary classification, and for filters of dimensions $n_{w}=\{31,7,11,11,3,3\}$, a dataset of 500 images and $n_{l}=2$ re-uploading layers, a perfect classification was achieved in all the cases, both in the training and testing sets.
@@ IV Conclusion
The quantum models were designed for a synthetic toy data set of phew grayscale pixels, considering the reduced number of available qubits and low connectivity in NISQ devices, and the computational cost of simulating them. An equivariant convolution operation has been proposed to implement the quantum circuits for higher-resolution images with color channels while preserving the label-invariance of the model. This way, the weights of convolution filters have been trained to learn the most essential features of the image, which are introduced as input to the quantum circuit. This hybrid algorithm has been tested on two public datasets to test its feasibility.
@@
Finally, we believe variational quantum algorithms have great potential as an early application of NISQ devices, although the anticipated quantum advantage may not come in the shape and flavor it is expected [60]. All in all, lots of work is to be done, and the geometric framework seems a powerful tool to design new architectures for machine learning models, even beyond quantum algorithms. We hope this work serves as an example of the methodology to build variational quantum classifiers with geometric inductive bias and encourages future research to propose new and innovative applications of quantum technologies.
Acknowledgements.- We acknowledge Ikerlan, DIPC, Ikerbasque, Basque Government, Diputación de Gipuzkoa and EIC for support, as well as discussions with the teams from Multiverse Computing and Ikerlan.

### v2

@@ III.5 Scaling the problem size
The data set employed in the previous sections represents a toy case created for the comparison of the capabilities of the geometric approach with respect to the traditional variational quantum models. The resolution of images in real computer vision problems, however, is far from the few pixels the current NISQ-era devices can handle due to their limited circuit size. A way to address this issue is to use a another encoding strategy, different from the one proposed in this work, that optimizes the number of features encoded per qubit. Another way to cope with this matter, rather than waiting to larger-scale quantum devices to be built and implement the proposed architecture for larger images, is to reduce the dimensionality of the input.
@@
where $r=(n_{f}-1)/2$ for odd $n_{f}$ and $r=n_{f}/2$ for even $n_{f}$, being $n_{f}$ the size of the filter. This expression can be extended for multichannel images, RGB images for example, by adding an extra dimension to the convolution filters, so the convolution filter acts over $n_{c}$ channels of the input tensor. However, in general, this operation is not equivariant to image rotations described by Eq. 9. For this reason, following the approach of Cohen, T., & Welling, M. [72], a rotation-equivariant convolution operation is defined by computing a convolution between an image and all rotated versions of the original filter, and averaging all resultant products into a single output (see Fig. 9).
@@
These equivariant convolutional layers can be stacked arbitrarily in $n_{c}^{\prime}$ channels, each one of them with a different filter so that they learn to focus on specific characteristics of the images. This way, if $n_{x}$ is the size (number of pixels) of the rows of the square image, $n_{w}$ the size of the filter, $s$ the stride and $p$ the padding, the size of the convoluted image will be [73]:
@@
The stride is the number of pixels the convolutional filter moves across the input image each time it slides and the padding is the addition of pixels (often zeros) around the input image’s border. For this case, $p=0$ and $s=1$ will be taken, so $n^{\prime}_{x}=n_{x}-n_{w}+1$. Taking the data set as a tensor of $N$ data-points allows us to apply the equivariant convolution operation over the whole batch, performing the following dimensional reduction:
@@
Accordingly, the number of channels of each convolutional layer and the size of the filters can be tuned to reduce the number of features and adapt the input to the size of the available quantum circuit. In this case, to connect the classical part to the equivariant quantum classifier described in the previous sections, the original images are reduced to $4\times 4$, single channel square images. It is postulated that while the number of qubits keeps increasing and the connectivity between them condensing, the size of the images that can be embedded directly into the quantum classifier will increase while decreasing the dependency on the classical convolutional part.
Therefore, the final model is constituted by a set of classical equivariant convolution layers, each one of them with optimizable parameters, that return an output in a reduced dimensional latent space, as an input for the rotation-invariant quantum model. The prediction of the model is evaluated in a mean quadratic distance loss function and the parameter update, both for the classical and quantum part, is done using Adam Optimizer to minimize the loss function, in the same fashion as in Sec. III.4. After each convolution layer, a ReLU activation function is applied in order to make the classical part learn nonlinearities [74]. Average pooling layers are applied between convolutions to reduce the dimensionality of the tensors. It can be verified that, unlike other types of pooling operations, average pooling is equivariant to image rotations as it is equivalent to performing a convolution with equal and normalized filter elements.
@@
This algorithm is tested on three public datasets. The first one, Fashion MNIST [^1] https://huggingface.co/datasets/fashion_mnist, contains $28\times 28$ grayscale images. For binary classification, labels 00 (T-shirts) and $1$ (trousers) have been taken. The second dataset, 11k Hands [^2] https://sites.google.com/view/11khands [75], contains $256\times 256$ RGB images of hands and the model is required to classify between left and right hands. It has been chosen as an illustrative example of a rotational symmetric problem where reflection symmetry is not present. Both previous datasets have been augmented by adding rotated versions of original images. The last dataset, Resisc45333https://huggingface.co/datasets/timm/resisc45 [76], contains satellital $256\times 256$ RGB images, and the model is required to classify between islands and lakes. All datasets are split with a test/train ratio of $1/3$. For reproducibility, the chosen hyperparameters for training the extended quantum models over each dataset are shown in Tab. 4.
Dataset $n_{w}$ $n_{c}$ $n_{p}$ $n_{l}$ Dataset Size Max. it. L.r.
MNIST 11,11,3,3 10,10,10,1 - 6 200 100 0.001
11kHands 3,2,2,4,7 8,64,64,6,1 2,2,2,2,2 6 100 1000 0.001
Resisc45 3,2,2,4,7 32,32,32,32,1 2,2,2,2,2 4 500 100 0.01
Table 4: Chosen hyperparameters for training the extended quantum models for each dataset. $n_{w}$ represents the dimension of each convolution filter, $n_{c}$ the number of channels in every convolutional layer, $n_{p}$ the dimension of each pooling filter, $n_{l}$ the number of re-uploading layers, Max. it. is the number of training epochs and L.r. is the learning rate of the optimizer.
Keeping the classical convolutional layers equal, different models with Equivariant, Non-equivariant and Basic-Entangler quantum circuits are trained over each dataset to compare the performances of the architectures. In Tab. 5 the performance metrics of each extended model are shown, for different datasets. It can be seen that the Equivariant models tend to outperform the other two architectures, pointing that the geometric approach does indeed bring a classification advantage where image rotation label symmetry is present. This is not true in the case of MNIST dataset, where Equivariant and Non-equivariant models show similar metrics. It is important to consider that classical convolution layers contribute significantly to the learning process, and there can be cases where the classical layers process the information enough to leave a trivial classification problem to the quantum part, vanishing any advantage coming from the geometric approach.
Dataset Model Type F1 Accuracy Precision Recall
MNIST Equiv. 0.91/0.89 0.89/0.86 0.89/0.87 0.88/0.86
MNIST Non-equiv. 0.92/0.88 0.90/0.86 0.90/0.87 0.90/0.96
MNIST Basic Entangler 0.83/0.81 0.77/0.75 0.76/0.76 0.77/0.75
11kHands Equiv. 0.84/0.40 0.85/0.57 0.83/0.43 0.86/0.57
11kHands Non-equiv. 0.53/0.24 0.71/0.60 0.60/0.38 0.71/0.60
11kHands Basic Entangler 0.57/0.28 0.78/0.62 0.65/0.36 0.78/0.62
Resisc45 Equiv. 0.64/0.64 0.78/0.80 0.59/0.88 0.78/0.80
Resisc45 Non-equiv. 0.52/0.52 0.70/0.71 0.72/0.86 0.70/0.71
Resisc45 Basic Entangler 0.38/0.44 0.64/0.68 0.53/0.88 0.64/0.68
Table 5: Performance metrics for each extended model type trained over different datasets. Each value in the table represents the mean performance metric, evaluated over the training set (left) and the test set (right). These metrics can be improved with additional computational resources, such as a deeper convolutional part, more re-uploading layers or a larger data set.
Finally, during the model training process, a loss landscape characterized by numerous local minima [77] was observed (see Fig. 10). From the study of the distribution of minima, we can extract some conclusions about the loss landscapes of each type of model. First, Tab. 6 shows that both the Equivariant and Non-equivariant models reach lower minima and they have a higher density in the minima below the mean loss value, comparing with the Basic Entangler model. According to Larocca M., et. al. [27], the required number of parameters to reach overparametrization is directly related to the dimension of the associated Dynamical Lie Algebra. In our case, all three models contain approximately the same number of parameters, but the dimension of the associated DLA is lower in the first two models. Therefore, the limit of overparametrization is closer for these two, and the trainability is expected to be improved compared to the Basic Entangler model. The observations agree with this principle. Appart from that, it can be seen how the lowest minima reached in every dataset correspond to the Equivariant model indicating that the introduced geometric bias helps the parameterized quantum circuit to produce better models. The study of the effect of initial points in the optimization process is considered of great interest and delegated to future work.
Dataset Model Type Mean Loss Min Loss % Below Mean
Tetrominoes Equiv $0.83$ $0.76$ 51
Tetrominoes Non-equiv $0.91$ $0.85$ $51$
Tetrominoes Basic Entangler $0.97$ $0.96$ $43$
MNIST Equiv $0.41$ $0.04$ $62$
MNIST Non-equiv $0.34$ $0.1$ $62$
MNIST Basic Entangler $0.7$ $0.27$ $48$
11kHands Equiv $0.6$ $0.24$ $65$
11kHands Non-equiv $0.71$ $0.34$ $70$
11kHands Basic Entangler $0.77$ $0.45$ $44$
Resisc45 Equiv $0.45$ $0.25$ $70$
Resisc45 Non-equiv $0.55$ $0.27$ $50$
Resisc45 Basic Entangler $0.81$ $0.68$ $40$
Table 6: Mean and minimum values of the loss function after the training of various randomly initialized parameters. The last column represents the percentage of loss observations below the mean value. Respectively, in order of appearance in the table, $35$, $21$, $9$ and $10$ random initializations were tried per model type in each dataset.
Figure 10: Cost history during the training of models. The black curve represents the mean cost in each step. Depending on the initialization of the parameters, the models get stuck in local minima. These results are for Resisc45 dataset, but the same phenomenon is observed for other datasets. 10 different initializations are shown for each model type.
@@ IV Conclusion
The quantum models were designed for a synthetic toy data set of phew grayscale pixels, considering the reduced number of available qubits and low connectivity in NISQ devices, and the computational cost of simulating them. An equivariant convolution operation has been proposed to implement the quantum circuits for higher-resolution images with color channels while preserving the label-invariance of the model. This way, the weights of convolution filters have been trained to learn the most essential features of the image, which are introduced as input to the quantum circuit. This hybrid algorithm has been tested on three public datasets and the three quantum architectures have been compared, obtaining again the best results for the geometric model. Numerical results have shown that the reduction of the associated Dynamical Lie Algebra, improves the trainability of the model by alleviating the problems of loss concentration (Barren Plateaus) and local minima, coinciding with theoretical results in literature. Gate symmetrization is a powerful tool for reducing the DLA while introducing inducted bias of the problem, improving its performance.
@@
Finally, we believe variational quantum algorithms have great potential as an early application of NISQ devices, although they still have significant limitations [78]. All in all, lots of work is to be done, and the geometric framework seems a powerful tool to design new architectures for machine learning models, even beyond quantum algorithms. We consider this work contributes in this direction, first, providing experimental evidence of the advantages of adding geometric inductive bias to the design of quantum circuits. Additionally, the employed methodology may serve as a blueprint for other research teams to apply GQML to image classification problems and push its state-of-the-art. We encourage researchers to use these results as a reference to benchmark and develop new applications. To conclude, we identify future research lines that extend this work, connecting theoretical concepts with experimental work (and vice versa), and interesting results are expected from them. We hope this work serves as a step in the path toward the utility of variational quantum algorithms.
Acknowledgements.- We acknowledge Ikerlan, DIPC, Ikerbasque, Basque Government, Diputación de Gipuzkoa and EIC for support, as well as discussions with the teams from Multiverse Computing and Ikerlan. This article has been supported by the Government of the Basque Country through the research grants ELKARTEK KUBIT - Grant number KK-2024/00105 (Kuantikaren Berrikuntzarako Ikasketa Teknologikoa), and the Gipuzkoa Provincial Council through the research grant QIA, grant number 2023-QUAN-000031-01 (Konputazio Kuantikoa Adimen Artifizialeko algoritmoetan: Konputazio klasikotik Quantum Machine Learning-era)